OpenStudy (anonymous):

PLEASE HELP!!!! just the answer please:) Solve the linear system. x + y = -3 y = 2x

OpenStudy (anonymous):

x=-1 y=-2

OpenStudy (anonymous):

what about, x + y = 6 x - y = 8 Solve the system of equations.

OpenStudy (anonymous):

x=7 y=-1

OpenStudy (anonymous):

if your not busy, 2x + y = 8 x + y = 4 The lines whose equations are given intersect at

OpenStudy (anonymous):

x=4 y=0

OpenStudy (anonymous):

3x + 2y = 5 5x + 2y = 7 Based on the given system of equations, which of the following is not true?

OpenStudy (anonymous):

x=1 y=1 I don't know what you mean by which is not true.

OpenStudy (anonymous):

oh, im sorry the options that I have are 2x=2 8x=12 4y=4

OpenStudy (anonymous):

The second one isn't true.

OpenStudy (anonymous):

would it be the first one?

OpenStudy (♪chibiterasu):

Please stop, you are not allowed to receive direct answers. It is against our code of conduct.

OpenStudy (anonymous):

k.thnks for the info

OpenStudy (anonymous):

do u just go around monitoring people?